The standard molar enthalpies of formation of cyclohexane(l) and benzene(l) at 25•c are -156 and 40 KJ/mol respectively.The standard enthalpy of hydrogenation of cyclohexene(l) at 25•c, is -119 KJ/mol.Find resonance energy of benzene: A.-152 kj/mol B.-159 kj/mol C.+152 kj/mol D.+159 Kj/mol

Grade:12th pass

1 Answers

Arkaprava Daripa
5 years ago
Enthalpy of formation of 3 C-C double bonds
= ΔHf(cyclohexane) - ΔHf(benzene)
= -156 - (49) kJ
= -205 kJ
Now it is given that,
Cyclohexene + H2 → cyclohexane ; ΔH = -119kJ
Now, theoretical enthalpy of formation of 3 C-C double bonds in benzene ring
=3*(-119)kJ
=-357kJ
So, resonance energy of benzene is = -357 - (-205)kJ
= -152 kJmol-1